OpenStudy (anonymous):

Simplify [ ( ½)^2 – ( ¼ )^3 ] x 2^3

OpenStudy (12345):

{ [(1/2)(1/2)]-[(1/4)(1/4)(1/4)] } x 2(2)(2) =[1/4-(1/16)(1/4)] x 8 =[ 8/4 - (8/16)(1/4) ] =2 - (1/2)(1/4) =2 - 1/8 =16/8 - 1/8 =15/8